# On arithmetic terms expressing the prime-counting function and the n-th prime

### v1

@@ Appendix A Monomials for the $n$-th prime equation $\hat{F}(n,\vec{x})=0$
$\begin{array}[]{ll}\hat{F}(n,\vec{x})=-2^{184x_{2}+288x_{3}+128x_{4}+35}x_{2}^{2}x_{23}^{3}x_{29}&+2^{328x_{2}+384x_{3}+128x_{4}+66}x_{2}^{2}x_{23}^{3}x_{29}\\ -2^{88x_{2}+224x_{3}+128x_{4}+11}x_{23}^{3}x_{29}x_{5}^{2}&-2^{328x_{2}+384x_{3}+128x_{4}+63}x_{2}x_{23}^{2}x_{28}\\ +2^{328x_{2}+384x_{3}+128x_{4}+64}x_{2}x_{23}^{3}x_{29}&-3\cdot 2^{32x_{2}+32x_{3}+40}x_{2}^{4}x_{22}^{5}x_{26}\\ -3\cdot 2^{40x_{2}+32x_{3}+46}x_{2}^{2}x_{22}^{5}x_{26}&-3\cdot 2^{32x_{2}+32x_{3}+39}x_{2}^{3}x_{22}^{5}x_{26}\\ +2^{88x_{2}+224x_{3}+128x_{4}+11}x_{23}^{2}x_{28}x_{5}&+3\cdot 2^{56x_{2}+32x_{3}+29}x_{2}^{3}x_{22}^{5}x_{26}\\ +3\cdot 2^{56x_{2}+32x_{3}+26}x_{2}^{2}x_{22}^{5}x_{26}&+3\cdot 2^{32x_{2}+32x_{3}+36}x_{2}^{2}x_{22}^{5}x_{26}\\ -2^{416x_{2}+608x_{3}+256x_{4}+76}x_{2}^{2}x_{5}^{2}&+2^{272x_{2}+512x_{3}+256x_{4}+45}x_{2}^{2}x_{5}^{2}\\ -3\cdot 2^{56x_{2}+32x_{3}+25}x_{2}x_{22}^{5}x_{26}&-2^{328x_{2}+384x_{3}+128x_{4}+61}x_{23}^{3}x_{29}\\ +3\cdot 2^{32x_{2}+32x_{3}+35}x_{2}x_{22}^{5}x_{26}&-2^{416x_{2}+608x_{3}+256x_{4}+74}x_{2}x_{5}^{2}\\ -1649267441647\cdot 2^{72x_{2}+32x_{3}+47}x_{2}^{2}&+1099511627787\cdot 2^{72x_{2}+32x_{3}+51}x_{2}^{4}\\ +1099511627809\cdot 2^{72x_{2}+32x_{3}+49}x_{2}^{3}&-2^{16x_{2}+32x_{3}+19}x_{2}^{4}x_{22}^{5}x_{26}\\ -2^{16x_{2}+32x_{3}+11}x_{22}^{5}x_{26}x_{5}^{4}&-3\cdot 2^{72x_{2}+64x_{3}+39}x_{2}^{3}x_{5}^{4}\\ -3\cdot 2^{72x_{2}+64x_{3}+36}x_{2}^{2}x_{5}^{4}&-3\cdot 2^{48x_{2}+64x_{3}+46}x_{2}^{2}x_{5}^{4}\\ -2^{32x_{2}+32x_{3}+35}x_{2}^{2}x_{22}^{3}x_{25}&-2^{16x_{2}+32x_{3}+11}x_{22}^{3}x_{25}x_{5}^{2}\\ -2^{280x_{2}+352x_{3}+128x_{4}+51}x_{23}x_{27}&+2^{40x_{2}+32x_{3}+51}x_{2}^{4}x_{22}^{5}x_{26}\\ +2^{56x_{2}+32x_{3}+31}x_{2}^{4}x_{22}^{5}x_{26}&+2^{40x_{2}+32x_{3}+49}x_{2}^{3}x_{22}^{5}x_{26}\\ +3\cdot 2^{48x_{2}+64x_{3}+50}x_{2}^{4}x_{5}^{4}&+3\cdot 2^{48x_{2}+64x_{3}+49}x_{2}^{3}x_{5}^{4}\\ +2^{56x_{2}+32x_{3}+26}x_{2}^{2}x_{22}^{3}x_{25}&-11\cdot 2^{56x_{2}+32x_{3}+21}x_{22}^{5}x_{26}\\ +3\cdot 2^{56x_{2}+64x_{3}+56}x_{2}^{2}x_{5}^{4}&-11\cdot 2^{32x_{2}+32x_{3}+31}x_{22}^{5}x_{26}\\ -2^{416x_{2}+608x_{3}+256x_{4}+73}x_{2}x_{5}&-2^{8x_{2}+18}x_{2}^{3}x_{24}x_{25}x_{28}x_{29}\\ -2^{8x_{2}+16}x_{2}^{2}x_{24}x_{25}x_{27}x_{29}&-2^{4x_{2}+11}x_{2}^{2}x_{24}x_{26}x_{27}x_{29}\\ +2^{284x_{2}+352x_{3}+128x_{4}+58}x_{2}x_{30}&+2^{280x_{2}+352x_{3}+128x_{4}+53}x_{2}x_{30}\\ -2^{512x_{2}+672x_{3}+256x_{4}+100}x_{2}^{4}&+2^{4x_{2}+14}x_{2}^{3}x_{24}x_{25}x_{28}x_{29}\\ +2^{4x_{2}+12}x_{2}^{2}x_{24}x_{25}x_{27}x_{29}&+2^{8x_{2}+15}x_{2}^{2}x_{24}x_{26}x_{27}x_{29}\\ +2199023255487\cdot 2^{72x_{2}+32x_{3}+44}x_{2}&-2^{512x_{2}+672x_{3}+256x_{4}+98}x_{2}^{3}\\ -2^{280x_{2}+352x_{3}+128x_{4}+55}x_{2}^{2}&+2^{656x_{2}+768x_{3}+256x_{4}+130}x_{2}^{4}\\ -2^{184x_{2}+288x_{3}+128x_{4}+31}x_{5}^{2}&+2^{656x_{2}+768x_{3}+256x_{4}+129}x_{2}^{3}\\ -3\cdot 2^{48x_{2}+64x_{3}+45}x_{2}x_{5}^{4}&+2^{656x_{2}+768x_{3}+256x_{4}+124}x_{2}^{2}\\ -2^{232x_{2}+320x_{3}+128x_{4}+45}x_{2}^{2}&-2^{136x_{2}+256x_{3}+128x_{4}+21}x_{5}^{2}\\ +2^{560x_{2}+704x_{3}+256x_{4}+104}x_{2}^{2}&+2^{368x_{2}+576x_{3}+256x_{4}+68}x_{2}^{4}\\ +2^{40x_{2}+32x_{3}+45}x_{2}x_{22}^{5}x_{26}&+2^{512x_{2}+672x_{3}+256x_{4}+95}x_{2}^{2}\\ +2^{424x_{2}+448x_{3}+128x_{4}+86}x_{2}^{2}&+2^{416x_{2}+608x_{3}+256x_{4}+71}x_{5}^{2}\\ +2^{176x_{2}+448x_{3}+256x_{4}+20}x_{5}^{4}&+3\cdot 2^{72x_{2}+64x_{3}+35}x_{2}x_{5}^{4}\\ +2^{56x_{2}+32x_{3}+24}x_{2}x_{22}^{3}x_{25}&-2^{4x_{2}+14}x_{2}^{3}x_{24}^{2}x_{25}x_{29}\\ +2^{376x_{2}+416x_{3}+128x_{4}+76}x_{2}^{2}&-2^{8x_{2}+17}x_{2}^{2}x_{24}x_{25}^{2}x_{28}\\ +2^{176x_{2}+448x_{3}+256x_{4}+20}x_{5}^{2}&+2^{8x_{2}+18}x_{2}^{3}x_{24}^{2}x_{25}x_{29}\\ +2^{4x_{2}+13}x_{2}^{2}x_{24}x_{25}^{2}x_{28}&-2^{4x_{2}+10}x_{2}x_{24}x_{25}x_{26}x_{27}\\ -2^{8x_{2}+14}x_{2}x_{25}x_{26}x_{27}x_{28}&-2^{656x_{2}+768x_{3}+256x_{4}+124}x_{2}\\ +2^{8x_{2}+14}x_{2}x_{24}x_{25}x_{26}x_{27}&+2^{4x_{2}+10}x_{2}x_{25}x_{26}x_{27}x_{28}\\ -2^{564x_{2}+704x_{3}+256x_{4}+108}x_{2}&-2^{288x_{2}+352x_{3}+128x_{4}+61}x_{30}\\ -2^{284x_{2}+352x_{3}+128x_{4}+56}x_{30}&-2^{72x_{2}+64x_{3}+41}x_{2}^{4}x_{5}^{4}\\ -2^{56x_{2}+64x_{3}+61}x_{2}^{4}x_{5}^{4}&-2^{56x_{2}+64x_{3}+59}x_{2}^{3}x_{5}^{4}\\ -2^{4x_{2}+14}x_{2}^{4}x_{24}^{2}x_{29}^{2}&-2^{72x_{2}+64x_{3}+36}x_{2}^{2}x_{5}^{2}\\ -21\cdot 2^{112x_{2}+64x_{3}+52}x_{2}^{4}&-2^{4x_{2}+12}x_{2}^{2}x_{24}^{2}x_{25}^{2}\\ -2^{4x_{2}+12}x_{2}^{2}x_{25}^{2}x_{28}^{2}&-275\cdot 2^{88x_{2}+64x_{3}+55}x_{2}^{2}\\ -25\cdot 2^{112x_{2}+64x_{3}+50}x_{2}^{3}&-119\cdot 2^{64x_{2}+64x_{3}+68}x_{2}^{4}\\ -2^{280x_{2}+352x_{3}+128x_{4}+53}x_{2}&-15\cdot 2^{88x_{2}+64x_{3}+68}x_{2}^{7}\\ -39\cdot 2^{88x_{2}+64x_{3}+65}x_{2}^{6}&+2^{32x_{2}+64x_{3}+29}x_{2}^{4}x_{5}^{4}\\ -21\cdot 2^{96x_{2}+64x_{3}+71}x_{2}^{4}&+2^{8x_{2}+18}x_{2}^{4}x_{24}^{2}x_{29}^{2}\\ +2^{424x_{2}+448x_{3}+128x_{4}+84}x_{2}&-37\cdot 2^{72x_{2}+64x_{3}+80}x_{2}^{4}\\ -33\cdot 2^{56x_{2}+32x_{3}+70}x_{2}^{4}&-13\cdot 2^{48x_{2}+32x_{3}+59}x_{2}^{4}\\ -33\cdot 2^{56x_{2}+32x_{3}+69}x_{2}^{3}&-33\cdot 2^{48x_{2}+32x_{3}+60}x_{2}^{4}\\ +2^{8x_{2}+16}x_{2}^{2}x_{24}^{2}x_{25}^{2}&+2^{8x_{2}+16}x_{2}^{2}x_{25}^{2}x_{28}^{2}\\ +2^{48x_{2}+64x_{3}+45}x_{2}^{2}x_{5}^{2}&-11\cdot 2^{40x_{2}+32x_{3}+41}x_{5}^{4}+\end{array}$
[…]

### v2

@@ Appendix A Monomials for the $n$-th prime equation $\hat{F}(n,\vec{x})=0$
$\begin{array}[]{ll}\hat{F}(n,\vec{x})=-2^{184x_{2}+288x_{3}+128x_{4}+35}x_{2}^{2}x_{23}^{3}x_{29}&+2^{328x_{2}+384x_{3}+128x_{4}+66}x_{2}^{2}x_{23}^{3}x_{29}\\ -2^{88x_{2}+224x_{3}+128x_{4}+11}x_{23}^{3}x_{29}x_{5}^{2}&-2^{328x_{2}+384x_{3}+128x_{4}+63}x_{2}x_{23}^{2}x_{28}\\ +2^{328x_{2}+384x_{3}+128x_{4}+64}x_{2}x_{23}^{3}x_{29}&-3\cdot 2^{32x_{2}+32x_{3}+40}x_{2}^{4}x_{22}^{5}x_{26}\\ -3\cdot 2^{40x_{2}+32x_{3}+46}x_{2}^{2}x_{22}^{5}x_{26}&-3\cdot 2^{32x_{2}+32x_{3}+39}x_{2}^{3}x_{22}^{5}x_{26}\\ +2^{88x_{2}+224x_{3}+128x_{4}+11}x_{23}^{2}x_{28}x_{5}&+3\cdot 2^{56x_{2}+32x_{3}+29}x_{2}^{3}x_{22}^{5}x_{26}\\ +3\cdot 2^{56x_{2}+32x_{3}+26}x_{2}^{2}x_{22}^{5}x_{26}&+3\cdot 2^{32x_{2}+32x_{3}+36}x_{2}^{2}x_{22}^{5}x_{26}\\ -2^{416x_{2}+608x_{3}+256x_{4}+76}x_{2}^{2}x_{5}^{2}&+2^{272x_{2}+512x_{3}+256x_{4}+45}x_{2}^{2}x_{5}^{2}\\ -3\cdot 2^{56x_{2}+32x_{3}+25}x_{2}x_{22}^{5}x_{26}&-2^{328x_{2}+384x_{3}+128x_{4}+61}x_{23}^{3}x_{29}\\ +3\cdot 2^{32x_{2}+32x_{3}+35}x_{2}x_{22}^{5}x_{26}&-2^{416x_{2}+608x_{3}+256x_{4}+74}x_{2}x_{5}^{2}\\ -1649267441647\cdot 2^{72x_{2}+32x_{3}+47}x_{2}^{2}&+1099511627787\cdot 2^{72x_{2}+32x_{3}+51}x_{2}^{4}\\ +1099511627809\cdot 2^{72x_{2}+32x_{3}+49}x_{2}^{3}&-2^{16x_{2}+32x_{3}+19}x_{2}^{4}x_{22}^{5}x_{26}\\ -2^{16x_{2}+32x_{3}+11}x_{22}^{5}x_{26}x_{5}^{4}&-3\cdot 2^{72x_{2}+64x_{3}+39}x_{2}^{3}x_{5}^{4}\\ -3\cdot 2^{72x_{2}+64x_{3}+36}x_{2}^{2}x_{5}^{4}&-3\cdot 2^{48x_{2}+64x_{3}+46}x_{2}^{2}x_{5}^{4}\\ -2^{32x_{2}+32x_{3}+35}x_{2}^{2}x_{22}^{3}x_{25}&-2^{16x_{2}+32x_{3}+11}x_{22}^{3}x_{25}x_{5}^{2}\\ -2^{280x_{2}+352x_{3}+128x_{4}+51}x_{23}x_{27}&+2^{40x_{2}+32x_{3}+51}x_{2}^{4}x_{22}^{5}x_{26}\\ +2^{56x_{2}+32x_{3}+31}x_{2}^{4}x_{22}^{5}x_{26}&+2^{40x_{2}+32x_{3}+49}x_{2}^{3}x_{22}^{5}x_{26}\\ +3\cdot 2^{48x_{2}+64x_{3}+50}x_{2}^{4}x_{5}^{4}&+3\cdot 2^{48x_{2}+64x_{3}+49}x_{2}^{3}x_{5}^{4}\\ +2^{56x_{2}+32x_{3}+26}x_{2}^{2}x_{22}^{3}x_{25}&-11\cdot 2^{56x_{2}+32x_{3}+21}x_{22}^{5}x_{26}\\ +3\cdot 2^{56x_{2}+64x_{3}+56}x_{2}^{2}x_{5}^{4}&-11\cdot 2^{32x_{2}+32x_{3}+31}x_{22}^{5}x_{26}\\ -2^{416x_{2}+608x_{3}+256x_{4}+73}x_{2}x_{5}&-2^{8x_{2}+18}x_{2}^{3}x_{24}x_{25}x_{28}x_{29}\\ -2^{8x_{2}+16}x_{2}^{2}x_{24}x_{25}x_{27}x_{29}&-2^{4x_{2}+11}x_{2}^{2}x_{24}x_{26}x_{27}x_{29}\\ -2^{512x_{2}+672x_{3}+256x_{4}+100}x_{2}^{4}&+2^{4x_{2}+14}x_{2}^{3}x_{24}x_{25}x_{28}x_{29}\\ +2^{4x_{2}+12}x_{2}^{2}x_{24}x_{25}x_{27}x_{29}&+2^{8x_{2}+15}x_{2}^{2}x_{24}x_{26}x_{27}x_{29}\\ +2199023255487\cdot 2^{72x_{2}+32x_{3}+44}x_{2}&-2^{512x_{2}+672x_{3}+256x_{4}+98}x_{2}^{3}\\ -2^{280x_{2}+352x_{3}+128x_{4}+55}x_{2}^{2}&+2^{656x_{2}+768x_{3}+256x_{4}+130}x_{2}^{4}\\ -2^{184x_{2}+288x_{3}+128x_{4}+31}x_{5}^{2}&+2^{656x_{2}+768x_{3}+256x_{4}+129}x_{2}^{3}\\ -3\cdot 2^{48x_{2}+64x_{3}+45}x_{2}x_{5}^{4}&+2^{656x_{2}+768x_{3}+256x_{4}+124}x_{2}^{2}\\ -2^{232x_{2}+320x_{3}+128x_{4}+45}x_{2}^{2}&-2^{136x_{2}+256x_{3}+128x_{4}+21}x_{5}^{2}\\ +2^{368x_{2}+576x_{3}+256x_{4}+68}x_{2}^{4}&+2^{40x_{2}+32x_{3}+45}x_{2}x_{22}^{5}x_{26}\\ +2^{512x_{2}+672x_{3}+256x_{4}+95}x_{2}^{2}&+2^{424x_{2}+448x_{3}+128x_{4}+86}x_{2}^{2}\\ +2^{416x_{2}+608x_{3}+256x_{4}+71}x_{5}^{2}&+2^{176x_{2}+448x_{3}+256x_{4}+20}x_{5}^{4}\\ +3\cdot 2^{72x_{2}+64x_{3}+35}x_{2}x_{5}^{4}&+2^{56x_{2}+32x_{3}+24}x_{2}x_{22}^{3}x_{25}\\ -2^{4x_{2}+14}x_{2}^{3}x_{24}^{2}x_{25}x_{29}&+2^{376x_{2}+416x_{3}+128x_{4}+76}x_{2}^{2}\\ -2^{8x_{2}+17}x_{2}^{2}x_{24}x_{25}^{2}x_{28}&+2^{176x_{2}+448x_{3}+256x_{4}+20}x_{5}^{2}\\ +2^{8x_{2}+18}x_{2}^{3}x_{24}^{2}x_{25}x_{29}&+2^{4x_{2}+13}x_{2}^{2}x_{24}x_{25}^{2}x_{28}\\ -2^{4x_{2}+10}x_{2}x_{24}x_{25}x_{26}x_{27}&-2^{8x_{2}+14}x_{2}x_{25}x_{26}x_{27}x_{28}\\ -2^{656x_{2}+768x_{3}+256x_{4}+124}x_{2}&+2^{8x_{2}+14}x_{2}x_{24}x_{25}x_{26}x_{27}\\ +2^{4x_{2}+10}x_{2}x_{25}x_{26}x_{27}x_{28}&-2^{288x_{2}+352x_{3}+128x_{4}+61}x_{30}\\ -2^{284x_{2}+352x_{3}+128x_{4}+56}x_{30}&-2^{72x_{2}+64x_{3}+41}x_{2}^{4}x_{5}^{4}\\ -2^{56x_{2}+64x_{3}+61}x_{2}^{4}x_{5}^{4}&-2^{56x_{2}+64x_{3}+59}x_{2}^{3}x_{5}^{4}\\ -2^{4x_{2}+14}x_{2}^{4}x_{24}^{2}x_{29}^{2}&-2^{72x_{2}+64x_{3}+36}x_{2}^{2}x_{5}^{2}\\ -21\cdot 2^{112x_{2}+64x_{3}+52}x_{2}^{4}&-2^{4x_{2}+12}x_{2}^{2}x_{24}^{2}x_{25}^{2}\\ -2^{4x_{2}+12}x_{2}^{2}x_{25}^{2}x_{28}^{2}&-275\cdot 2^{88x_{2}+64x_{3}+55}x_{2}^{2}\\ -25\cdot 2^{112x_{2}+64x_{3}+50}x_{2}^{3}&-119\cdot 2^{64x_{2}+64x_{3}+68}x_{2}^{4}\\ -15\cdot 2^{88x_{2}+64x_{3}+68}x_{2}^{7}&-39\cdot 2^{88x_{2}+64x_{3}+65}x_{2}^{6}\\ +2^{32x_{2}+64x_{3}+29}x_{2}^{4}x_{5}^{4}&-21\cdot 2^{96x_{2}+64x_{3}+71}x_{2}^{4}\\ +2^{8x_{2}+18}x_{2}^{4}x_{24}^{2}x_{29}^{2}&+2^{424x_{2}+448x_{3}+128x_{4}+84}x_{2}\\ -37\cdot 2^{72x_{2}+64x_{3}+80}x_{2}^{4}&-33\cdot 2^{56x_{2}+32x_{3}+70}x_{2}^{4}\\ -13\cdot 2^{48x_{2}+32x_{3}+59}x_{2}^{4}&-33\cdot 2^{56x_{2}+32x_{3}+69}x_{2}^{3}\\ -33\cdot 2^{48x_{2}+32x_{3}+60}x_{2}^{4}&+2^{8x_{2}+16}x_{2}^{2}x_{24}^{2}x_{25}^{2}\\ +2^{8x_{2}+16}x_{2}^{2}x_{25}^{2}x_{28}^{2}&+2^{48x_{2}+64x_{3}+45}x_{2}^{2}x_{5}^{2}\\ -11\cdot 2^{40x_{2}+32x_{3}+41}x_{5}^{4}&+2^{376x_{2}+416x_{3}+128x_{4}+74}x_{2}\\ -15\cdot 2^{64x_{2}+32x_{3}+77}x_{2}^{2}&-33\cdot 2^{48x_{2}+32x_{3}+59}x_{2}^{3}\\ -11\cdot 2^{32x_{2}+32x_{3}+39}x_{2}^{4}&-2^{8x_{2}+15}x_{2}x_{24}x_{25}^{2}x_{27}+\end{array}$
[…]